# The X-ray rise and fall of the Recurrent Symbiotic System T CrB

### v1

@@ 4 Discussion
Within the assumptions of the analysis presented here, the proposed disk model can reproduce satisfactorily the different epochs of observations. The best-fit is achieved with an inclination angle between the symmetry axis of the disk and the line of sight of $\theta=50^{\circ}$, which is very close to the value of 60$\pm$5 reported by Belczynski & Mikolajewska (1998). However, it is very likely that a more realistic accretion disk model, with an intricate 3D spiral structure as that found by hydrodynamical simulations (e.g., Liu et al., 2017), is needed to reproduce the fine details of the observations (a situation that is also suggested by the CF parameter). For example, the current model is not sufficient to completely reproduce the emission in the 1.0–4.0 keV energy range for the 2018.16 and 2022.27 epochs (see Fig. 4 and 5). Still, the analysis presented here demonstrates that the contribution from the reflection component should not be easily discarded as it significantly contributes to the total X-ray flux for some of the epochs and helps fitting the Fe complex and dominates for energies above 15 keV. Furthermore, in some cases, the reflection component has a larger contributing flux ($F_{\mathrm{X,Ref}}$) than other components (see Table 2 and the top left panel of Fig. 6).
@@
In general, the analysis of the X-ray spectra from the 2006.77+2015.81, 2017.24 and 2018.16 epochs resulted in properties that are very similar to those reported by previous authors (Luna et al., 2008, 2018; Kennea et al., 2009; Zhekov & Tomov, 2019). The heavily extinguished plasma component $kT_{2}$, very likely produced at the boundary layer, has a larger temperature for the steady-state epoch (2006.77+2015.81) and then is reduced after the start of the active phase. The rise of the soft component is coincident with the beginning of the active phase. Nevertheless, as previously stated, the rise of the soft component is not always coupled with a decay of the hard one.
@@
The temperature of the heavily-extinguished plasma ($kT_{2}$) is estimated to be 14.8$\pm$1.6 keV in the steady-state epoch (2006.77$+$2015.81), with properties very similar to those estimated by previous authors (Luna et al., 2008; Kennea et al., 2009). After the increase in the optical activity in 2016.35, the analysis of the X-ray observations suggests that $kT_{2}$ has decreased, going to 2.8$\pm$0.8 keV in the 2017.24 epoch to an averaged value of $\approx$8.0 keV in the more recent years (2018.16–2022.66). Changes in the temperature of the boundary layer can be used to assess $\dot{M}_{\mathrm{acc}}$ onto the WD component. Theoretical estimations have shown that high plasma temperatures are expected for low $\dot{M}_{\mathrm{acc}}$ and vice versa (e.g., Pringle & Savonije, 1979; Patterson & Raymond, 1985). In particular, Patterson & Raymond (1985) used a simple accretion disk model and found that low $\dot{M}_{\mathrm{acc}}$ makes the boundary layer to have low density and low optical depth, which does not cool efficiently and remains at high temperatures. On the other hand, high $\dot{M}_{\mathrm{acc}}$ produce a denser boundary layer that becomes optically thicker, cooling down and radiating its energy at lower temperatures. These authors proposed that $\dot{M}_{\mathrm{acc}}$ and the plasma temperature of the boundary layer (in our case $kT_{2}$) are related through the equation
@@
with $M_{\mathrm{WD}}$ as the mass of the WD. Estimated to be 1.2 M in the case of T CrB (Belczynski & Mikolajewska, 1998).
Using this equation, we illustrate the variation of $\dot{M}_{\mathrm{acc}}$ as a function of time in Fig. 6 (top right panel). This plot results in an expected behaviour, where $\dot{M}_{\mathrm{acc}}$ is lower in the steady-state phase (2006.77+2015.81 epoch) with a value of 9.6×$\times 10^{-11}$ M yr-1 and then increases after the start of the active phase. $\dot{M}_{\mathrm{acc}}$ has values of 5.1×$\times 10^{-10}$, 2.4×$\times 10^{-10}$, 1.9×$\times 10^{-10}$, 1.5×$\times 10^{-10}$, 1.9×$\times 10^{-10}$ and 1.6×$\times 10^{-10}$ M yr-1 for the 2017.24, 2018.16, 2020.77, 2021.73, 2022.27 and 2022.66 epochs, respectively. Previous $\dot{M}_{\mathrm{acc}}$ estimations for T CrB resulted in higher values, but we attribute these to the fact that other works assume that the flux of the soft component also corresponds to accreted material.
Other authors have estimated $\dot{M}_{\mathrm{acc}}$ by using the formula
$L_{\mathrm{acc}}=\frac{GM_{\mathrm{WD}}}{R_{\mathrm{WD}}}\dot{M}_{\mathrm{acc}},$ (3)
were $G$ as the gravitational constant, $R_{\mathrm{WD}}$ as the radius of the WD component and $L_{\mathrm{acc}}$ is the luminosity produced by the accretion process. Typically, most works just assume that the X-ray luminosity is the pure signature of accretion, and simply use the X-ray luminosity to estimate $\dot{M}_{\mathrm{acc}}$. As an exercise, we use $L_{\mathrm{X2}}$ in Eq. (3) and also plot the results in the top right panel of Fig. 6 alongside those obtained from Eq. (2). This formula suggests that the mass accretion has reduced in the latest years which is somehow contradictory to the evident increase in activity revealed by the optical light curves of T CrB (see, e.g., Fig. 2).
@@
The increase in the accretion rate of material onto the WD is somehow inversely proportional to the total X-ray flux. There is a clear anti-correlation between the total X-ray flux $F_{\mathrm{X,TOT}}$ and the optical measurements. Fig. 6 shows that $F_{\mathrm{X,TOT}}$ was higher by the 2006.77 epoch and experienced a reduction already shown by 2017.24. A minimum in $F_{\mathrm{X,TOT}}$ was reached between the 2018.16 and 2020.77 epochs and started recovering in the later epochs of observation.
@@
A reduction of the ratio between the X-ray emission and that of the optical band ($F_{\mathrm{X}}/F_{\mathrm{opt}}$) is expected for larger $\dot{M}_{\mathrm{acc}}$ (see, e.g., fig. 7 of Patterson & Raymond, 1985). To corroborate this, averaged optical fluxes are estimated from the $V$- and $B$-magnitude measurements by averaging half a year around the same dates as the X-ray observations. These averaged fluxes, calculated for each epoch are also listed at the bottom of Table 2. Fig. 6 bottom left panel shows the evolution with time of $F_{\mathrm{X}}/F_{\mathrm{opt}}$ for the $V$ and $B$ bands, while the bottom right panel shows the dependence of this ratio with respect to $\dot{M}_{\mathrm{acc}}$. In general, the trend is the reduction of $F_{\mathrm{X}}/F_{\mathrm{opt}}$, except for the 2017.24 epoch, which is illustrated by the rightmost points in the bottom right panel of Fig. 6, the later dominated by the strong soft X-ray emission produced by the jet.
@@
Fig. 7 shows the HID of T CrB where we compare the total X-ray flux $F_{\mathrm{X,TOT}}$ versus the observed hardness ratio defined as the ratio of the 3.0–10.0 keV band over the 0.3–3.0 keV band. T CrB has evolved similarly to that described by, for example, Motta et al. (2021) for black hole X-ray binaries. The evolution starts and finishes at the bottom rightmost part of the diagram (Low Hard) where, evidently, we do not have any X-ray observation, but is the region where the latest observations are converging to. This is denoted by a (blue) star in Fig. 7. It seems like, after the nova-like event experienced during 1946.1, T CrB entered the steady-state phase represented by the 2006.77 epoch going into the High Hard region of the HID. An evolutionary path that is expected to repeat after the next nova-like event predicted to occur during September of the present year. Very likely after the outburst experienced in 2016.35, this symbiotic system moved to the Soft part of the diagram illustrated by the 2017.24 epoch. The subsequent evolution moved T CrB into the Low part of the Soft and Hard intermediate phases. Vertical dashed lines are shown in the figure for illustration purposes separating the 2018.16, 2020.77 and 2022.27 epochs, where the soft X-ray component is still considerably, from those of 2021.73 and 2022.66 where the soft X-ray emission has almost vanished (see Fig. 4 and 5). We note that T CrB went back and forth between the Soft Intermediate and Hard Intermediate phases between 2021.73 and 2022.66, a characteristic that is not unusual in other compact accreting systems (see Motta et al., 2021, and references therein).
T CrB is practically the first symbiotic system with such detailed monitoring that allowed us to trace an almost complete loop in the HID, going from the steady-state phase into just a couple of years before its next nova-like event. This opens a new window for the studies of accreting WDs in symbiotic systems, adding to the multi-scale reach of accretion physics around compact objects. It would be interesting to corroborate that such evolution in the HID is the same for all symbiotic systems, however, there are not many targets with sufficient time coverage.
@@ 5 Summary
We presented the analysis of archival X-ray observations of the recurrent symbiotic star T CrB obtained with NuSTAR, Suzaku and XMM-Newton covering the 2006.77–2022.66 period. The analysis of the X-ray spectra was performed by including a reflection model produced by the presence of the accretion disk. The combination of these data in addition to optical monitoring allowed us to study the variable behaviour of T CrB in a time span of about 16 years. Our main findings can be summarised as follows:
@@
  • The spectral analysis suggests that a disk structure with a radius of $R=100$ AU, a width of $h$=10 AU, averaged column density $N_{\mathrm{H,Ref}}=10^{25}$ cm-2 and an inclination of $\theta$=50 between the light of sight and its symmetry axis can explain all of the X-ray spectra, naturally reproducing the 6.4 keV Fe emission line. The reflection component is necessary to explain the emission in the 10.0–50.0 keV spectral range detected in the NuSTAR data, dominating the $E>15$ keV energy range. Although previous authors argued that the reflection component should dominate the 2.0–4.0 keV energy range, in T CrB there seems to be a competition between the properties of the plasma temperature of the boundary layer ($kT_{2}$ in our model) and the reflection component.
  • The presence of emission lines in the 2017.24 XMM-Newton data prevents from adopting a black body emission model to fit the soft spectral range. Instead, we use optically-thin emission plasma models that resulted in an average temperature of $kT_{1}$=3.5×$\times 10^{5}$ K. Although these temperatures are also observed in $\alpha$-type symbiotic systems their luminosities are orders of magnitude larger. We then suggest that these plasma temperatures are best explained by the presence of jet-like emissions with velocity values ranging from 110 to 200 km s-1, an explanation already invoked to explain the softest emission in other $\beta/\delta$ symbiotic systems.
  • As expected from theoretical predictions, the evolution of the plasma temperature of the boundary layer ($kT_{2}$) is consistent with the increase of the activity in T CrB. A high plasma temperature of 14.8 keV is observed in epochs before the onset of the current active phase which resulted in an estimated mass accretion rate of $\dot{M}_{\mathrm{acc}}$=9.6×$\times 10^{-11}$ M yr-1 using the predictions from a simple accretion disk from Patterson & Raymond (1985). Conversely, lower plasma temperatures in the current active phase ($kT_{2}$=2.8–9.5 keV) translate into mass accretion rates of $\dot{M}_{\mathrm{acc}}$=[1.5–5.1]×$\times 10^{-10}$ M yr-1.
@@
T CrB is a unique case of a recurrent symbiotic systems for several reasons. It is one of the few symbiotic stars with available multi-epoch, multi-instrument X-ray observations that have observed the steady-state phase and its evolution through the active phase. T CrB is of particular interest given that it is predicted that by September 2024 it will produce its next nova-like outburst.
@@ Acknowledgements
J.A.T. and D.A.V.T. acknowledge support from the UNAM PAPIIT project IN102324. O.G.M. thanks Fundación Marcos Moshinsky (Mexico), UNAM PAPIIT project IN109123 and CONAHCyT "Frontera de la Ciencia" project CF-2023-G100. JAT thanks Emilio Tejeda for helping with the sketch of the disk model presented here and for long discussion sessions that helped draft the first version of this manuscript. D.A.V.T. thanks Consejo Nacional de Humanidades, Cientica y Tecnología (CONAHCyT, Mexico) for a student grant. This work is based on observations obtained with XMM-Newton, an European Science Agency (ESA) science mission with instruments and contributions directly funded by ESA Member States and NASA. This research also made use of data obtained from the Suzaku satellite, a collaborative mission between the space agencies of Japan (JAXA) and the USA (NASA). This research has made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work has made extensive use of NASA’s Astrophysics Data System.

### v2

@@ 4.1 The disk properties
In fact, individual epochs might likely require different disk structures with slightly different parameters as demonstrated by Zamanov et al. (2024). This can be particularly seen for the 2018.16 and 2022.27 epochs which have some difficulties fitting the 0.9–3.0 keV energy range. However, a more realistic accretion disk model, with an intricate 3D spiral structure as that found by hydrodynamical simulations (e.g., de Val-Borro et al., 2017; Liu et al., 2017; Lee et al., 2022), might be needed to reproduce the fine details of the observations (a situation that is also suggested by the CF parameter). Still, the analysis presented here demonstrates that the contribution from the reflection component should not be easily discarded as it significantly contributes to the total X-ray flux for some of the epochs and helps fitting the Fe complex and dominates the spectrum for energies above 15 keV. Furthermore, in some cases, the reflection component has a larger contributing flux ($F_{\mathrm{X,Ref}}$) than other components (see Table 2 and the top panel of Fig. 6).
4.2 On the presence of a bipolar ejection
@@
In general, the analysis of the X-ray spectra from the 2006.77+2015.81, 2017.24 and 2018.16 epochs resulted in properties that are very similar to those reported by previous authors (Luna et al., 2008, 2018; Kennea et al., 2009; Zhekov & Tomov, 2019). The heavily extinguished ($N_{\mathrm{H2}}>10^{23}$ cm-2) component $kT_{2}$, very likely produced at the boundary layer, has a larger temperature for the steady-state epoch (2006.77+2015.81) and then is reduced after the start of the active phase. The rise of the soft component is coincident with the beginning of the active phase. Nevertheless, as previously stated, the rise of the soft component is not always coupled with a decay of the hard one.
@@
4.3 Consequences for the temporal evolution of T CrB
The temperature of the heavily-extinguished plasma ($kT_{2}$) is estimated to be 14.8$\pm$1.6 keV in the steady-state epoch (2006.77$+$2015.81), with properties very similar to those estimated by previous authors (Luna et al., 2008; Kennea et al., 2009). After the increase in the optical activity in 2016.35, the analysis of the X-ray observations suggests that $kT_{2}$ has decreased, going to 2.8$\pm$0.8 keV in the 2017.24 epoch to an averaged value of $\approx$8.0 keV in the more recent years (2018.16–2022.66).
Changes in the temperature of the boundary layer can be used to assess $\dot{M}_{\mathrm{acc}}$ onto the WD component. Theoretical estimations have shown that high plasma temperatures are expected for low $\dot{M}_{\mathrm{acc}}$ and vice versa (e.g., Pringle & Savonije, 1979; Patterson & Raymond, 1985). In particular, Patterson & Raymond (1985) used a simple accretion disk model and found that low $\dot{M}_{\mathrm{acc}}$ makes the boundary layer to have low density and low optical depth, which does not cool efficiently and remains at high temperatures. On the other hand, high $\dot{M}_{\mathrm{acc}}$ produce a denser boundary layer that becomes optically thicker, cooling down and radiating its energy at lower temperatures. Patterson & Raymond (1985) proposed that $\dot{M}_{\mathrm{acc}}$ and the plasma temperature of the boundary layer (in our case $kT_{2}$) are related through the equation
@@ 4.3 Consequences for the temporal evolution of T CrB
with $M_{\mathrm{WD}}$ as the mass of the WD. Estimated to be 1.32 M in the case of T CrB (Shara et al., 2018).
Using Eq. (2) we estimate that $\dot{M}_{\mathrm{acc}}$ is lower in the steady-state phase (2006.77+2015.81 epoch) with a value of 1.4×$\times 10^{-10}$ M yr-1 and then increases after the start of the active phase to an averaged value of 3.4×$\times 10^{-10}$ for the 2017.24–2022.66 epochs. However, we note that previous $\dot{M}_{\mathrm{acc}}$ estimations for T CrB resulted in higher values, particularly reaching to values $\gtrsim 10^{-8}$ M yr-1 (see, e.g., table 1 in Zamanov et al., 2023).
On the other hand, we can estimate $\dot{M}_{\mathrm{acc}}$ by assuming that the accretion process produces a certain bolometric luminosity $L_{\mathrm{acc}}$ such as
$L_{\mathrm{acc}}=\frac{GM_{\mathrm{WD}}}{\cos\theta~{}R_{\mathrm{WD}}}\frac{\dot{M}_{\mathrm{acc}}}{2},$ (3)
were $G$ as the gravitational constant and $R_{\mathrm{WD}}$ is the radius of the WD component. By assuming that the signature of accretion is the X-ray luminosity of the boundary layer (in our case that corresponding to $F_{\mathrm{X2}}$), we obtained $M_{\mathrm{acc}}$ values of a few times $10^{-9}$, where the largest value is that of the steady state phase and still over an order of magnitude below previous predictions.
To be able to use Eq. (3) we need to account for all the luminosity produced by accretion in the different bands. As a lower limit, we can approximate $L_{\mathrm{acc}}$ only by accounting the luminosities obtained in X-rays and the optical bands. That is, by simply assuming $L_{\mathrm{acc}}\approx L_{\mathrm{X}}+L_{\mathrm{opt}}$. Taking advantage of the $L_{\mathrm{opt}}$ estimations of the accretion disk presented in Zamanov et al. (2023) for the active phase of T CrB we estimate $\dot{M}_{\mathrm{acc}}$ values of 2.4×$\times 10^{-8}$, 4.1×$\times 10^{-8}$, 2.0×$\times 10^{-8}$, 2.4×$\times 10^{-8}$, 2.3×$\times 10^{-8}$ and 2.4×$\times 10^{-8}$ for the 2017.24, 2018.16, 2020.77, 2021.73, 2022.72 and 2022.66 epochs. These $\dot{M}_{\mathrm{acc}}$ estimates are also listed in the bottom row of Table 2 and are further illustrated in the middle panel of Fig. 6. It is evidently expected that the mass accretion rate to be smaller for the pre-active phase.
Given that the X-ray luminosity is about $\approx$1–10 per cent of $L_{\mathrm{opt}}$, our $\dot{M}_{\mathrm{acc}}$ estimations only represent slightly improved values than those estimated by Zamanov et al. (2023). We particularly note the case of the 2017.24 epoch that has contemporary optical observations in Zamanov et al. (2023). These authors estimate $\dot{M}_{\mathrm{acc}}$=1.8×$\times 10^{-8}$ M yr-1 and accounting for the X-ray emission this is estimated to be about 30 per cent larger. These results illustrate that applying Eq. (3) by only adopting the X-ray luminosity of the boundary layer one would need to apply efficiency factors $\eta\approx$20–100 to recover more realistic $\dot{M}_{\mathrm{acc}}$ values.
@@
The increase in the accretion rate of material onto the WD is somehow inversely proportional to the total X-ray flux. There is a clear anti-correlation between the total X-ray flux $F_{\mathrm{X,TOT}}$ and the optical measurements. The top panel of Fig. 6 shows that $F_{\mathrm{X,TOT}}$ was higher by the 2006.77 epoch and experienced a reduction already shown by 2017.24. A minimum in $F_{\mathrm{X,TOT}}$ was reached between the 2018.16 and 2020.77 epochs and started recovering in the later epochs of observation.
A reduction of the ratio between the X-ray emission and that of the optical band ($F_{\mathrm{X}}/F_{\mathrm{opt}}$) is expected for larger $\dot{M}_{\mathrm{acc}}$ (see, e.g., fig. 7 of Patterson & Raymond, 1985). Fig. 6 bottom panel shows the evolution with time of $F_{\mathrm{X}}/F_{\mathrm{opt}}$ for the $B$, $V$ and $I$ bands which can be directly compared with the evolution of the $\dot{M}_{\mathrm{acc}}$ in the middle panel. To calculate optical fluxes and luminosities from the AAVSO $I$, $V$ and $B$ photometric observations, we estimated median values for time intervals of half a year around the same dates as the X-ray observations. These fluxes ($<F_{I}>$, $<F_{V}>$ and $<F_{B}>$), calculated for each epoch are also listed at the bottom of Table 2.
@@
Fig. 7 shows the HID of T CrB where we compare the total X-ray flux $F_{\mathrm{X,TOT}}$ versus the observed hardness ratio defined as the ratio of the 3.0–10.0 keV band over the 0.3–3.0 keV band. T CrB has evolved similarly to that described by, for example, Motta et al. (2021) for black hole X-ray binaries. The evolution starts and finishes at the bottom rightmost part of the diagram (Low Hard) where, evidently, we do not have any X-ray observation, but is the region where the latest observations are converging to. This is denoted by a (blue) star in Fig. 7. It seems like, after the nova-like event experienced during 1946.1, T CrB entered the steady-state phase represented by the 2006.77 epoch going into the High Hard region of the HID. An evolutionary path that is expected to repeat after the next nova-like event predicted to occur in the near future. Very likely after the outburst experienced in 2016.35, this symbiotic system moved to the Soft part of the diagram illustrated by the 2017.24 epoch. The subsequent evolution moved T CrB into the Low part of the Soft and Hard intermediate phases. Vertical dashed lines are shown in the figure for illustration purposes separating the 2018.16, 2020.77 and 2022.27 epochs, where the soft X-ray component is still considerably, from those of 2021.73 and 2022.66 where the soft X-ray emission has almost vanished (see Fig. 4 and 5). We note that T CrB went back and forth between the Soft Intermediate and Hard Intermediate phases between 2021.73 and 2022.66, a characteristic that is not unusual in other compact accreting systems (see Motta et al., 2021, and references therein).
T CrB is practically the first symbiotic system with such detailed monitoring that allowed us to trace an almost complete loop in the HID, going from the steady-state phase into just a couple of years before its next nova-like event. This opens a new window for the studies of accreting WDs in symbiotic systems, adding to the multi-scale reach of accretion physics around compact objects. It would be interesting to corroborate that such evolution in the HID is the same for other symbiotic recurrent nova systems (for example, RS Oph) or any other accreting WD in a binary system.
@@ 5 Summary
We presented the analysis of archival X-ray observations of the symbiotic recurrent nova system T CrB obtained with NuSTAR, Suzaku and XMM-Newton covering the 2006.77–2022.66 period. The analysis of the X-ray spectra was performed by including a reflection model produced by the presence of the accretion disk. The combination of these data in addition to optical monitoring allowed us to study the variable behaviour of T CrB in a time span of about 16 years. Our main findings can be summarised as follows:
@@
  • The spectral analysis suggests that a disk structure with a radius of $R=$1 AU, a width of $h$=0.1 AU, averaged column density $N_{\mathrm{H,Ref}}=10^{25}$ cm-2 and an inclination of $\theta$=50 between the light of sight and its symmetry axis can explain all of the X-ray spectra, naturally reproducing the 6.4 keV Fe emission line. The reflection component is necessary to explain the emission in the 10.0–50.0 keV spectral range detected in the NuSTAR data, dominating the $E>15$ keV energy range. The proposed reflecting disk structure is of the same order of the Roche lobe radius estimated for the WD component, only a factor of $\sim$2 larger than the sizes estimated by Zamanov et al. (2024). Our disk model seem to suggest that not only the accretion disk produces reflection, but some of this effect might be also produced by the high-density material in the vicinity of the accretion disk.
  • The presence of emission lines in the 2017.24 XMM-Newton data prevents from adopting a black body emission model to fit the soft spectral range. Instead, we use optically-thin emission plasma models that resulted in an average temperature of $kT_{1}$=3.5×$\times 10^{5}$ K. Although these temperatures are also observed in $\alpha$-type symbiotic systems their luminosities are orders of magnitude larger. We then suggest that these plasma temperatures are best explained by the presence of variable jet-like emissions with velocity values ranging from 110 to 200 km s-1, an explanation already invoked to explain the softest emission in other $\beta/\delta$ symbiotic systems (e.g., R Aqr, HM Sge and V694 Mon).
  • As expected from theoretical predictions, the evolution of the plasma temperature of the boundary layer ($kT_{2}$) is consistent with the increase of the activity in T CrB. A high plasma temperature of 14.8 keV is observed in epochs before the onset of the current active phase. Conversely, lower plasma temperatures in the current active phase ($kT_{2}$=2.8–9.5 keV).
  • Using optical and X-ray flux estimates, we slightly improved previous the estimations for the mass accretion rate $\dot{M}_{\mathrm{acc}}$. An averaged value of 2.6×$\times 10^{-8}$ M yr-1 for the current active phase.
@@
T CrB is a unique case of a symbiotic recurrent nova system for several reasons. It is one of the few symbiotic stars with available multi-epoch, multi-instrument X-ray observations that have observed the steady-state phase and its evolution through the active phase. T CrB is of particular interest given that it is predicted to experience its next nova-like outburst in the near future.
@@ Acknowledgements
The authors are in debt to an anonymous referee for comments and suggestions that helped improve the presentation and discussion of the results. J.A.T. and D.A.V.T. acknowledge support from the UNAM PAPIIT project IN102324. O.G.M. thanks Fundación Marcos Moshinsky (Mexico), UNAM PAPIIT project IN109123 and CONAHCyT "Frontera de la Ciencia" project CF-2023-G100. JAT thanks Emilio Tejeda for helping with the sketch of the disk model presented here and for long discussion sessions that helped draft the first version of this manuscript. D.A.V.T. thanks Consejo Nacional de Humanidades, Cientica y Tecnología (CONAHCyT, Mexico) for a student grant. This work is based on observations obtained with XMM-Newton, an European Science Agency (ESA) science mission with instruments and contributions directly funded by ESA Member States and NASA. This research also made use of data obtained from the Suzaku satellite, a collaborative mission between the space agencies of Japan (JAXA) and the USA (NASA). This research has made use of data from the NuSTAR mission, a project led by the California Institute of Technology, managed by the Jet Propulsion Laboratory, and funded by the National Aeronautics and Space Administration. We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research. This work has made extensive use of NASA’s Astrophysics Data System.
@@
Appendix A Temporal evolution in optical bands
Fig. 8 presents the temporal evolution of the optical light curves of T CrB obtained thought the $I$, $V$ and $B$ filters obtained from AAVSO.
Figure 8: Light curve of T CrB obtained from photometric measurements obtained with the $I$, $V$ and $B$ filters from the AAVSO. The vertical lines represent the epochs in which XMM-Newton observations are available and analysed in this paper.